OpenStudy (anonymous):

A pizza with a 14in diameter (7in radius) is cut into 10 slices. Another pizza with the same diameter and radius is cut into 12 slices. Find the interior angle and arc length for each pizza, as well as the area for each slice.

OpenStudy (anonymous):

@ParthKohli or @Miss_Deva Either of you know how to do this?><

OpenStudy (anonymous):

Even if you only show me how to do one of the pizzas so I can do the second

OpenStudy (callisto):

For pizza cut into 10 slices angle of each slice = 360/10 = 36 \[arc \ length \ of \ each \ slice = 2 \pi r \times \frac{\theta}{360}=2 \pi (7) \times \frac{36}{360}\]\[Area \ of \ each \ slice = \pi (7)^2 \times \frac{36}{360}\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (callisto):

welcome!~

OpenStudy (anonymous):

@Callisto so do I just leave them at 2π(7)×36/360 and π(7)2×36/360 Or do I solve them all the way down to the exact number?

OpenStudy (callisto):

at least you need to simplify it!~ leave in terms of pi or decimal/integer, depends on the requirement of the question

OpenStudy (anonymous):

Alrighty! Thank you!

OpenStudy (anonymous):

@Callisto Sorry for bugging you on this again! But for the arc length AND the area I got an estimated 4.396 Are they supposed to be the same?

OpenStudy (callisto):

Area you've got is wrong :| 7^2 = 7 x 7 = 49...

OpenStudy (anonymous):

I wasnt squaring it>< my badd Geometry is not my cup of tea

OpenStudy (callisto):

Not my cup of tea too :|